If (Mr. X + Mrs. Y = 69 years) and (Mr. X + Mr. Z = 72 years) and (Mrs. Y + Mr. Z = 79 years).
So, how old are Mr. X, Mrs. Y and Mr. Z?

Found 2 solutions by ikleyn, MathTherapy:
Answer by ikleyn(52802)

.
If (Mr. X + Mrs. Y = 69 years) and (Mr. X + Mr. Z = 72 years) and (Mrs. Y + Mr. Z = 79 years).
So, how old are Mr. X, Mrs. Y and Mr. Z?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

X + Y     = 69,      (1)
X     + Z = 72,      (2)
    Y + Z = 79.      (3)

Add the equations (1), (2) and (3). You will get

2X + 2Y + 2Z = 69 + 72 + 79 = 220,  

or

X + Y + Z = 110.     (4)

Now distract (1) from (4). You will get 
Z = 110 - 69 = 41.

Now distract (2) from (4). You will get 
Y = 110 - 72 = 38.

Now distract (3) from (4). You will get 
X = 110 - 79 = 31.

Answer. X = 31, Y = 38, Z = 41.

For similar problems, see the lesson The tricks to solve some word problems with three and more unknowns using mental math in this site.


Answer by MathTherapy(10552)


If (Mr. X + Mrs. Y = 69 years) and (Mr. X + Mr. Z = 72 years) and (Mrs. Y + Mr. Z = 79 years).
So, how old are Mr. X, Mrs. Y and Mr. Z?